# ICONS: Influence Consensus for Vision-Language Data Selection

### v3

@@ 3.3 Analysis of aggregation strategies
Table 3: Comparison of aggregation approaches. Performance of different influence aggregation methods when selecting 20% of the LLAVA-665K dataset. Our proposed aggregation approach (Vote) consistently achieves the best overall performance (98.6% Rel.), outperforming both score-based (Merge, Max), their noramlized variants (Merge-GausNorm, Merge-SumNorm) and rank-based (Round Robin, MinRank) baselines.
Limitations of score-based aggregation.  Score-based methods like Merge (Eqn. 3) and Max (Eqn. 4) assume calibrated influence scores across tasks, which is rarely the case. We observe substantial variation in the distribution of influence scores across tasks with standard deviations spanning from $8.15\times 10^{-3}$ (MME) to $1.26\times 10^{-3}$ (VQAv2), indicating that influence scores for MME are much more spread out, while those for VQAv2 are tightly concentrated. Similarly, mean influence scores vary in both magnitude and sign: MME has a relatively high positive mean ($1.68\times 10^{-3}$), while tasks like POPE ($-2.83\times 10^{-4}$) and GQA ($-8.89\times 10^{-5}$) have negative means. These divergent patterns show that certain tasks have much wider influence scores distributions, making a sample helpful for one task but neutral or harmful for another. Aggregating raw scores thus biases selection toward tasks with higher variance or skewed means. To address this calibration issue, we experimented with normalization strategies: Merge-SumNorm, which rescales each task’s influence scores by dividing them by a task-specific normalization factor (i.e., sum), and Merge-GausNorm, which normalizes the scores using task-wise mean and standard deviation before averaging:
$I_{\text{Merge-SumNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})}{\sum_{j}I_{k}({\bm{z}}_{j})}\qquad I_{\text{Merge-GausNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
Limitations of rank-based aggregation.  Rank-based methods sidesteps the challenge of comparing raw influence scores by focusing on within-task ranking. Round Robin selects samples by cycling through each task and picking the highest-scoring remaining sample for that task, ensuring balanced coverage. MinRank (Eqn. 5) selects samples that have the best minimum rank across all tasks, prioritizing examples that perform exceptionally well in at least one task regardless of their performance in others. Although these methods ensure balanced coverage across tasks, they can overfit to outlier tasks. This is particularly evident with LLaVA-W Bench llava , which is an outlier in its influence ranking: both Round Robin and MinRank achieve relatively high scores on it (e.g. MinRank: 68.4), but this comes at the cost of lower performance on all other tasks (Tab. 3). This suggests that purely rank-based selection can trade off some overall efficacy on the mainstream tasks and hurt multi-task balance. In contrast, our Vote approach is more robust and avoids this by focusing on multi-task consensus rather than forcing equal representation, yielding better balance and higher overall performance (98.6% Rel.), highlighting the importance of identifying broadly influential examples rather than narrowly optimizing per-task rankings.
Divergent multi-task influence patterns.  As shown in Fig. 4, the pairwise overlap heatmap shows notable variation in training data influence across tasks. High overlap – e.g., VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks. However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others. Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples. To understand task-specific influence matrices from the specialist stage, we select the top 20% samples per task (Specialists). Overlap with our generalist subset (Fig. 4) varies significantly, from minimal in tasks like VQAv2 (3.27%) and VizWiz (3.28%) to substantial agreement in tasks like LLAVA-W Bench llava (24.21%). These findings empirically demonstrate significant overlap in influential samples across tasks and validate our approach: by analyzing task-specific gradient-based influence patterns and building consensus across tasks, we can identify a compact subset that captures broadly useful samples across tasks, yielding strong performance with significantly less data.
3.4 ICONS generalizes to unseen tasks
LLaVA-ICONS-133K demonstrates exceptional generalization on entirely unseen benchmarks that were not used during data selection. As shown in Tab. 4, we test across a diverse spectrum of tasks including MMVet mmvet , NaturalBench naturalbench , AI2D ai2d , ChartQA chartqa , DocVQA docvqa , InfoVQA infographicvqa , RealWorldQA xai2024grok and CMMMU cmmmu . LLaVA-ICONS-133K achieves 95.5-113.9% (Rel.) compared to full dataset training (InfoVQA: 103.8%, NaturalBench: 105.5%, RealWorldQA: 104.4%, and CMMMU: 113.9%), suggesting that, in some cases, training on LLaVA-ICONS-133K may even outperform training on the full dataset, despite these tasks not being included in the selection process. Importantly, LLaVA-ICONS-133K significantly outperforms random selection across all benchmarks. This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains. We further provide the cross-architecture generalization results in Appendix §D.
Full 55.4 17.5 28.9 26.5 31.1 12.4 52.4 22.1 100.0
Random 50.2 15.1 25.2 24.3 27.6 11.1 49.8 21.9 91.6
LLaVA-ICONS-133K 53.9 17.1 27.9 27.5 29.7 12.8 55.0 25.2 98.7
Per-task Rel. (%) 97.3 97.7 96.5 103.8 95.5 103.2 104.4 114.0 -
Table 4: Unseen-task generalization. Performance comparison on unseen benchmarks when trained on selected subsets. Notably, we observe improvements on InfoVQA (103.8%), RealWorldQA (104.4%), and CMMMU (114.0%), highlighting strong generalization to unseen tasks.
Figure 5: Different selection ratios. ICONS consistently outperforms all baselines across different selection ratios and remarkably exceeding 102% at 60% selection ratio.
To understand how ICONS scales, we evaluate it across different selection ratios, progressively scaling the subset size from 5% to 60% of LLaVA-665K. As shown in Fig. 5, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial. Second, as the selection ratio increases, the performance gap between different methods gradually narrows. This convergence pattern is expected, as larger sample sizes naturally capture more of the dataset’s diversity and information. Despite this convergence trend, ICONS consistently outperforms all baselines across all selection ratios. Remarkably, it even surpasses full dataset performance at the 60% ratio, achieving over 102% relative score. One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
4 Conclusion
In this work, we introduce ICONS, a simple yet effective influence consensus-based approach for visual instruction tuning data selection. By leveraging gradient-based influence estimation and aggregating task-specific selections through majority voting, our two-stage specialist-to-generalist approach selects training examples that are broadly beneficial across multiple downstream tasks. ICONS addresses limitations of prior selection methods by avoiding assumptions about score comparability across tasks and reducing the sensitivity to outlier task rankings, which can bias selection in both score-based and rank-based approaches. Beyond data selection, it provides a principled way to reason about data influence in multitask mixtures. Through extensive experiments, we show that ICONS builds compact, high-impact datasets without sacrificing performance or generalization, achieving 98.6% of full dataset performance using only 20% of LLaVA-665K and generalizing well to unseen tasks or architectures. We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, 20% subsets of the LLaVA-665K and Cambrian-7M datasets, maintaining strong performance on diverse tasks and transferring well to unseen ones. We hope our work inspires further exploration into data-efficient methods for vision-language models across diverse applications.
Appendix A Computational complexity
A.1 Complexity analysis
A.2 Resource requirements
@@ A.3 Discussion on cost-benefit justification
Appendix B Related work
B.1 Visual instruction tuning
Multimodal large language models (MLLMs), e.g., Flamingo flamingo , LLaVA llava , BLIP2 blip-2 , and Cambrian (tong2024cambrian, ), enhance the capabilities of large language models (LLMs) on various multimodal tasks. A key component in advancing MLLMs is visual instruction tuning (llava, ), a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants. This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications (llava, ).
B.2 Data selection
Data selection methods survey can be categorized based on the types of information they utilize for selection. Representation-based approaches (semdedup, ; coincide, ) leverage neural embeddings to capture data representations. Loss trajectory-based methods (mindermann2022prioritized, ) prioritize data points that contribute most significantly to reducing generalization error over training. Gradient-based techniques (el2n, ; xia2024less, ; deng2024influential, ) select data based on gradient information. Recent work has explored various approaches to select optimal visual instruction tuning datasets. Concurrent work TIVE (liu2024less, ) employs gradient-based selection to identify representative instances. TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty. Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks. COINCIDE (coincide, ) clusters data based on representations associated with concept-skill compositions. Our work follows targeted instruction tuning selection approach LESS xia2024less to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information from various tasks and selecting data for multiple downstream tasks via majority voting.
Appendix C Additional experiment details & ablations
Task MME POPE SQA-I MMBench VQAv2 GQA VizWiz TextVQA LLaVA-W
@@ Appendix C Additional experiment details & ablations
$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164 1,000 398 800 84 84
$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784 36,807 12,578 8,000 5,000 84
Task Type Y/N Y/N MCQ MCQ MCQ VQA VQA VQA VQA VQA
C.1 Additional task details
C.2 Projection dimension
We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions. The choice of 5120 was empirically validated for its trade-off between effective capturing gradient representation and maintaining a manageable parameter space. Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count. We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig. 6(a).
C.3 Warm-up Ratio
@@
Appendix D Additional analysis
In this section, we provide additional analyses to better understand different aspects of our ICONS. We begin by analyzing the effectiveness of task-specific selections and their aggregation into a generalist subset (§D.1). We further explore whether incorporating visual dependency information into the selection process affects performance across different types of vision-language tasks (§D.2). We also evaluate the transferability of our selected subset across different model scales (§D.3). Additionally, we evaluate the consistency of our method across multiple runs (§D.4), showing its robustness and reliability.
@@ Appendix D Additional analysis
Delta (%) 1.04 0.65 5.65 -1.43 0.18 -1.04 1.35 4.34 1.06 1.49
D.1 From specialist to generalist
@@
D.2 Visual dependency influence ranking
Recent work tong2024cambrian has shown that vision-language tasks vary in their reliance on visual information: tasks like MMBench mmbench depend heavily on visual grounding, while others like SQA-I sqa can be handled primarily through language, showing only a 5% drop in performance when visual input is removed tong2024cambrian . To take visual dependency of training data into account, we further explored gradient-based Visual Dependency Score (VDS). For each data point, we calculate the gradient of the model’s auto-regressive cross-entropy loss with both the original image and a Gaussian noise image $I_{\text{noise}}\sim\mathcal{N}(0,1)$, keeping the text input constant. This quantifies how much the visual component contributes to model performance. We construct an adapted influence matrix: visual influence matrix $\mathcal{I}_{\texttt{VDS}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}|}$, which quantifies the visual influence of each training sample ${\bm{z}}_{i}$ on each validation sample ${\bm{z}}^{\prime}_{j}$ w.r.t the model’s gradient alignment and visual dependency. $\mathcal{I}_{\text{VDS}}$ is computed as:
Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different vision-language tasks (Tab. 7). The approach shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%). However, tasks that primarily rely on textual reasoning show decreased performance, including SQA-I (-1.84%) and TextVQA (-1.44%). These results align with and extend the findings in Cambrian tong2024cambrian , demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly. This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in training.
@@ D.2 Visual dependency influence ranking
D.3 Cross-architecture generalization
We further conduct experiments on cross architecture generalization to evaluate the transferability of our selected data across different model scales. While our subset was initially selected using LLaVA-v1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-v1.5-13B. This tests whether our selection criteria identify universally valuable training examples rather than model-specific patterns. Our results in Tab. 8 show cross-architecture generalization, with 13B model trained on 7B-selected subset achieving 98.1% Rel.. Both 7B-selected and 13B-selected subsets outperform random selection (95.7%), with the 13B-selected option showing particular strength in reasoning tasks like MMBench and POPE. This suggests that our selected subset captures fundamental visual-language understanding patterns that generalize well across different model architectures.
@@ D.3 Cross-architecture generalization
D.4 Consistency analysis
To evaluate the consistency of ICONS, we conduct three independent runs of the experiments. As shown in Fig. 7, our method demonstrates high consistency across different runs, achieving 98.6$\pm$1.2Rel., which shows a notable improvement over random baseline, which achieves 95.8$\pm$2.7%. The lower standard deviation (1.2% vs 2.7%) further indicates that our approach produces more stable and reliable outcomes compared to the random baseline.
Appendix E Algorithm details
@@
14:$S\leftarrow\text{top-}p\text{ samples by }I_{\text{vote}}$
@@
Appendix F Limitations
Our approach primarily faces one practical limitation: computing gradients for large training datasets is computationally expensive (Appendix §A). This computational overhead could potentially constrain the method’s applicability when working with extremely large-scale datasets. To support broader research community, we release LLaVA-ICONS-133K dataset to help research iteration and model development under resource-constrained settings.
Appendix G Broader impact
Our exploration focuses on scientific understanding and practical applications of vision-language data selection. While our work does not directly imply negative impacts, it may indirectly propagate existing biases present in the original datasets. Therefore, it is important to incorporate rigorous bias-mitigation measurements for data selection. On the positive side, our method enables more efficient and sustainable model development by reducing data redundancy, computational cost while maintaining or even improving performance. Discussion on these critical aspects should remain a priority as we further explore the potential of vision-language data selection.
Appendix H Future work
@@
Appendix I Visualizations
I.1 Representation-based vs. Gradient-based data selection
While ICONS leverages gradient-based influence signals, we explore how representation-based data selection (RDS) performs in the same setting (§3.2). We analyze the top-ranked training examples selected by RDS after the generalist stage in Fig. 8 vs. samples selected by ICONS in Fig. 19. Interestingly, we observe that the representation-based variants often favor training examples with repeated images or instructions, which may dominate the learned representations without contributing to better generalization. Some of the highest-scoring samples under representation-based similarity are duplicated image-question pairs with only the answer choices shuffled. We hypothesize that this is a side effect of the way multimodal representations are constructed—where visually dominant or textually redundant samples occupy high-density regions in embedding space. However, these samples do not necessarily translate into broader utility across tasks, as seen in the performance gap in Tab. 1. This discrepancy raises broader questions about what it means for multimodal data to be diverse. While we leave these questions open for future exploration, our results suggest that gradient-based influence, though computationally more expensive, is better aligned with generalization and multi-task data mixture settings.

### v4

@@ A.3 Discussion on cost-benefit justification
Task MME POPE SQA-I MMBench
@@
$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164
$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784
[…]
Task VQAv2 GQA VizWiz TextVQA LLaVA-W
$|\mathcal{D}_{\text{val}}|$ 1,000 398 800 84 84
$|\mathcal{D}_{\text{test}}|$ 36,807 12,578 8,000 5,000 84
[…]
Appendix B Additional experiment details & ablations
We provide comprehensive experimental details and ablation studies to support our main findings. We first present detailed task descriptions and dataset statistics (Tab. 6), followed by detailed per-task results for aggregation method comparisons (Tab. 5) and full results for unseen benchmarks (Tab. 7). We then ablate key hyperparameters including projection dimension (§B.2) and warm-up ratio (§B.3).
Table 7: Detailed results of unseen task generalization. Detailed results for Fig. 3. Performance comparison on unseen benchmarks when trained on selected 20% subsets. Notably, we observe improvements on InfoVQA (103.8%), RealWorldQA (104.4%), and CMMMU (114.0%), highlighting strong generalization to unseen tasks.
Full 55.4 17.5 28.9 26.5 31.1 12.4 52.4 22.1 100.0
Random 50.2 15.1 25.2 24.3 29.6 11.1 49.8 21.9 91.8
CLIP-Score 52.0 16.4 27.1 24.9 29.2 11.6 49.3 20.8 93.9
EL2N 52.1 17.9 26.9 26.4 29.6 12.5 51.9 21.7 97.8
Perplexity 53.0 16.4 27.1 27.1 29.4 11.4 50.3 23.9 97.0
SemDeDup 53.2 16.4 26.4 26.6 29.1 13.2 51.1 22.8 97.8
D2-Pruning 52.1 17.3 26.5 24.7 33.1 11.3 52.0 21.9 96.7
Self-Sup 52.7 16.6 27.4 25.2 29.6 11.8 50.1 21.0 95.0
Self-Filter 52.1 16.0 27.1 25.0 29.3 11.5 49.5 20.6 94.0
COINCIDE 53.7 17.0 27.9 26.0 30.2 12.1 50.9 21.4 97.2
RDS 53.0 16.8 27.5 25.4 29.8 12.0 50.3 21.3 95.6
LLaVA-ICONS-133K 53.9 17.1 27.9 27.5 29.7 12.8 55.0 25.2 101.6
[…]
@@ Appendix B Additional experiment details & ablations
B.1 Additional task details
B.2 Projection dimension
We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions. The choice of 5120 was empirically validated for its trade-off between effectively capturing gradient representations and maintaining a manageable parameter space. Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count. We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig. 5(a).
B.3 Warm-up Ratio
@@
[…]
Appendix C Additional analysis
In this section, we provide additional analyses to better understand different aspects of our ICONS. We include detailed vote distribution visualizations (Fig. 6) extending the analysis from §4.6. We analyze the effectiveness of task-specific selections and their aggregation into a generalist subset (§C.1). We further explore whether incorporating visual dependency information into the selection process affects performance across different types of vision-language tasks (§C.2). We also evaluate the transferability of our selected subset across different model scales (§C.3). Additionally, we evaluate the consistency of our method across multiple runs (§C.4), showing its robustness and reliability.
Figure 6: Vote distributions across selection ratios. Each line shows the vote distribution for a different selection ratio (5% to 95%). At lower ratios (e.g., 5%), most samples receive zero votes, resulting in a left-skewed distribution. As ratios increase (e.g., 50%, 70%, 90%), the distribution shifts rightward and flattens, with more samples receiving votes from multiple tasks. Stricter criteria (lower ratios) identify task-specific samples, while relaxed criteria (higher ratios) reveal broader multi-task agreement.
C.1 From specialist to generalist
@@
Divergent multi-task influence patterns.  As shown in Fig. 7, the pairwise overlap heatmap shows notable variation in training data influence across tasks. High overlap – e.g., VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks. However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others. Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples. To understand task-specific influence matrices from the specialist stage, we select the top 20% samples per task (Specialists). Overlap with our generalist subset (Fig. 8) varies significantly, from minimal in tasks like VQAv2 (3.27%) and VizWiz (3.28%) to substantial agreement in tasks like LLAVA-W Bench [llava] (24.21%). These findings empirically demonstrate significant overlap in influential samples across tasks and validate our approach: by analyzing task-specific gradient-based influence patterns and building consensus across tasks, we can identify a compact subset that captures broadly useful samples across tasks, yielding strong performance with significantly less data.
@@
C.2 Visual dependency influence ranking
Recent work [tong2024cambrian] has shown that vision-language tasks vary in their reliance on visual information: tasks like MMBench [mmbench] depend heavily on visual grounding, while others like SQA-I [sqa] can be handled primarily through language, showing only a 5% drop in performance when visual input is removed [tong2024cambrian]. To take visual dependency of training data into account, we further explored gradient-based Visual Dependency Score (VDS). For each data point, we calculate the gradient of the model’s auto-regressive cross-entropy loss with both the original image and a Gaussian noise image $I_{\text{noise}}\sim\mathcal{N}(0,1)$, keeping the text input constant. This quantifies how much the visual component contributes to model performance. We construct an adapted influence matrix: visual influence matrix $\mathcal{I}_{\texttt{VDS}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}|}$, which quantifies the visual influence of each training sample ${\bm{z}}_{i}$ on each validation sample ${\bm{z}}^{\prime}_{j}$ w.r.t the model’s gradient alignment and visual dependency. $\mathcal{I}_{\text{VDS}}$ is computed as:
Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different tasks (Tab. 9). It shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%). However, tasks that primarily rely on textual reasoning show decreased performance, including SQA-I (-1.84%) and TextVQA (-1.44%). These results align with and extend the findings in Cambrian [tong2024cambrian], demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly. This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in training.
@@ C.2 Visual dependency influence ranking
C.3 Cross-architecture generalization
We conduct experiments on cross architecture generalization to evaluate the transferability of our selected data across different model scales. While our subset was initially selected using LLaVA-v1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-v1.5-13B. This tests whether our selection criteria identify universally valuable training examples rather than model-specific patterns. Our results in Tab. 10 show cross-architecture generalization, with 13B model trained on 7B-selected subset achieving 98.1% Rel.. Both 7B-selected and 13B-selected subsets outperform random selection (95.7%), with the 13B-selected option showing particular strength in reasoning tasks like MMBench and POPE. This suggests our selected subset captures fundamental visual-language understanding patterns that generalize well across different model architectures.
@@ C.3 Cross-architecture generalization
C.4 Consistency analysis
[…]
Appendix D Algorithm details
@@
[…]
@@
Appendix E Future work
@@
Appendix F Visualizations
In this section, we provide detailed visualizations. We compare representation-based versus gradient-based data selection approaches (§F.1), showing qualitative differences in selected samples, and visualize the most influential examples identified by both task-specific specialists and the generalist model across all target tasks (§F.2). Additionally, we present examples of low-vote training samples (Fig. 10), which are referenced in §4.7 of the main paper, to illustrate the characteristics of data filtered out by our method.
F.1 Representation-based vs. Gradient-based data selection
While ICONS leverages gradient-based influence signals, we explore how representation-based data selection (RDS) performs in the same setting (§4.2). We compare the top-ranked training samples selected after the generalist stage by RDS with those selected by ICONS in Fig. 22. Interestingly, we observe that the representation-based variants often favor training examples with repeated images or instructions, which may dominate the learned representations without contributing to better generalization. Some of the highest-scoring samples under representation-based similarity are duplicated image-question pairs with only the answer choices shuffled. We hypothesize that this is a side effect of the way multimodal representations are constructed, where visually dominant or textually redundant samples occupy high-density regions in embedding space. However, these samples do not necessarily translate into broader utility across tasks, as seen in the performance gap in Tab. 1. This discrepancy raises broader questions about what it means for multimodal data to be diverse. While we leave these questions open for future exploration, our results suggest that gradient-based influence, though computationally more expensive, is better aligned with generalization and multi-task data mixture settings.
Figure 10: Examples of 0 & 1-vote training samples. These low-vote samples exhibit characteristics that explain their low selection frequency: (1) subjective questions lacking objective ground truth (e.g., “Is the woman pretty?”), (2) trivial questions providing minimal visual learning signal (e.g., “Is the door open?”), (3) redundant questions about the same image content. By filtering out such potentially harmful or redundant examples, ICONS focuses training on more informative and generalizable samples.